OpenStudy (anonymous):

Find the x-intercept(s), if any exist, of the parabola defined by the function below. f(x) = (4x + 12)(x + 4) A. (48,0) B. no x-intercept(s) C. (-3,0) and (-4,0) D. (3,0) and (4,0)

OpenStudy (nirmalnema):

(-3,0) and (-4,0)

OpenStudy (haseeb96):

just multiply them f(x) = (4x+12)(x+4) = 4x^2 + 16x +12x +48 = 4x^2 +28x +48 = 4 ( x^2 + 7x +12) 4(x^2+7x+12)=0 x^2+7x+12=0 now apply the factorizing method x^2 +3x+4x+12=0 x(x+3) +4(x+3)=0 (x+4)(x+3)=0 either x+4=0 and x+3=0 x=-4 and x=-3 as u know that on x intercept we take y=0 so the points are (-3,0) and (-4,0) this is the correct answer

OpenStudy (haseeb96):

i hope u would get it

OpenStudy (haseeb96):

@rosebird Read it

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@Haseeb96